# Solving Dense Linear Systems Faster than via Preconditioning

### v1

@@ 1 Introduction
Solving a linear system $\mathbf{A}\mathbf{x}=\mathbf{b}$ is one of the most classical algorithmic tasks, with applications in scientific computing, engineering, data science, optimization, and more. Many classical deterministic algorithms are known for this task, from direct methods such as Gaussian elimination to iterative methods such as Conjugate Gradient. More recently, there has been much success in developing faster algorithms based on randomized techniques such as sampling and sketching for solving linear systems under specific structural conditions, e.g., graph-structured [ST14, KMP12, ${}^{+}$18] and other classes of sparse linear systems [XXG12, Gra06], as well as extremely over/under-determined systems [Sar06, Woo14]. However, the benefits of using randomized algorithms for general dense linear systems remain unclear.
@@
The above discussion indicates that solving linear systems with a matrix having only $k=o(n)$ large singular values should be substantially easier than in the worst-case. Understanding this phenomenon is motivated by the real-world prevalence of such linear systems: either due to direct regularization (e.g., $\mathbf{A}=\mathbf{B}+\lambda\mathbf{I}$), as in continuous optimization [BV04], machine learning [ZDW13], and statistics [DW18], or because the input matrix is distorted by isotropic noise (e.g., $\mathbf{A}=\mathbf{B}+\delta\mathbf{N}$, where $\mathbf{N}$ is the noise), which can be caused by measurement error [LW11], data compression [${}^{+}$21], and privacy enforcement [DR14]. Yet, the $\tilde{O}(n^{2}k)$ complexity of CG is highly unsatisfactory, as it only provides an improvement over direct methods for small $k$, and does not lead to a near-linear time algorithm unless $k=\tilde{O}(1)$. This leads to our central question:
@@
Arguably the most popular strategy of accelerating the complexity of linear system solvers such as CG is via preconditioning. For instance, given access to $\mathbf{A}$’s best rank $k$ approximation $\mathbf{A}_{k}$, we can easily construct a preconditioner matrix $\mathbf{M}$ that allows CG to solve our problem in $O(\log(1/\epsilon))$ iterations (since $\mathbf{A}_{k}$ describes the entire ill-conditioned part of $\mathbf{A}$’s spectrum). While computing $\mathbf{A}_{k}$ exactly is as expensive as directly solving the linear system, much recent work has been dedicated to fast randomized algorithms for low-rank approximation via matrix sketching. In particular, we can use power iteration [HMT11] (or Krylov subspace iteration [MM15]) to construct a low-rank approximation $\tilde{\mathbf{A}}_{k}$ with a spectral norm guarantee $\|\tilde{\mathbf{A}}_{k}-\mathbf{A}\|_{2}=O(\sigma_{k+1}(\mathbf{A}))$, sufficient to construct a good preconditioner for our problem. These algorithms require (repeatedly) multiplying matrix $\mathbf{A}$ with an $n\times k$ random matrix, leading to the runtime of $\tilde{O}(n^{2}k^{\omega-2})$. This complexity can be further refined using fast rectangular matrix multiplication [LG12] by replacing the exponent $\omega$ with $\omega(1,1,\log_{n}(k))\leq\omega$. Nevertheless, for large enough $k$, the resulting preconditioning algorithm will not run in near-linear time. [^1] Using current matrix multiplication algorithms, we have $\omega(1,1,\alpha)>2$ for any $\alpha>0.321335$ [WXXZ23]. For simplicity, from now on we will present all time complexities using just the square matrix multiplication exponent $\omega$. Some randomized low-rank approximation algorithms avoid power iteration and run in $\tilde{O}(n^{2}+nk^{\omega-1})$ time [CW13, ${}^{+}$15, CMM17], but they only guarantee a bound on the Frobenius norm error $\|\tilde{\mathbf{A}}_{k}-\mathbf{A}\|_{F}$ which is worse by up to a $\sqrt{n}$ factor than the spectral norm bound and thus does not in general lead to a good preconditioner.
@@ 1.2 Our Techniques
A core inspiration for our algorithms is a broad family of stochastic iterative methods for solving linear systems, which includes randomized Kaczmarz [SV09], coordinate descent [LL10], sketch-and-project [GR15], and many others [LS13]. Unlike CG-type methods which do a pass over the entire matrix $\mathbf{A}$ in each iteration, these approaches randomly sub-sample a small number of rows/columns of $\mathbf{A}$ (or construct a small sketch of $\mathbf{A}$) and use this stochastic information to perform a sublinear iterative update. There has been much success in characterizing the convergence and complexity of stochastic methods in the ultra-small sample regime (i.e., single row/column per iteration [SV09]), but the theoretical analysis of the larger sample regime has proven much more elusive. While there has been some progress in this direction over the last few years, it has come with trade-offs: meaningful characterizations of the convergence behavior have required relying on either constructing sophisticated and expensive sub-sampling distributions (such as determinantal point processes [RK20, MDK20]), or even more expensive dense sketching methods (such as the Gaussian sketch [RN21, DR22]), rendering the entire procedure computationally impractical.
@@ 1.3 Background and Related Work
There has also been much work on accelerating iterative linear system solvers, particularly those based on Krylov subspace methods, such as conjugate gradient / Lanczos. Here, most of the efforts have been concentrated on constructing effective preconditioners, for example via specialized efficient Cholesky factorizations [${}^{+}$16, KS16], or on improving their numerical stability [PV21, Nie22]. However, these approaches have focused on sparse and/or structured systems, e.g., those decomposable into a sum of simple components, particularly Laplacians of undirected graphs [Vai89, ST14, KMP12], and their extensions [${}^{+}$18]. Other examples of structured systems with efficient solvers include Hankel/Toeplitz matrices [KKM79, XXG12] and circulant matrices [Gra06].
@@
Stochastic solvers. The literature on stochastic iterative methods for solving linear systems has its origins in the Kaczmarz method [Kac37], which performs iterative updates while cycling through the rows of the system and projecting the current iterate onto the hyperplane formed by an individual row. A randomized version of this method was introduced and analyzed by Strohmer and Vershynin [SV09], followed up by closely related randomized coordinate descent methods for psd linear systems and least squares by Leventhal and Lewis [LL10]. This triggered much research into stochastic iterative solvers, leading to more general classes of methods that sample blocks of rows instead of a single row [NW13, NT14, MDK20], or use matrix sketching [RN21, DR22] via the Sketch-and-Project framework [GR15]. Despite these efforts, it is still not fully understood when stochastic solvers are faster than deterministic (e.g., Krylov-based) iterative methods. One such result was obtained by Lee and Sidford [LS13], who showed that the convergence rate of an accelerated stochastic coordinate descent solver exhibits better condition number dependence than conjugate gradient for psd linear systems. However, this still leads to poor performance for matrices that have an ill-conditioned top-$k$ part of the spectrum, whereas our algorithms avoid this by using blocks of coordinates (rows or columns) as opposed to one coordinate at a time.
@@
Matrix sketching and sub-sampling. Our techniques are closely related to matrix sketching and sampling techniques commonly used in randomized linear algebra [HMT11, Woo14, DM16, MT20, ${}^{+}$23], which is an area that has led to improved randomized approximation algorithms for tasks such as least squares regression [Sar06, RT08, CW13], low-rank approximation [${}^{+}$15, CMM17, LW20], $l_{p}$ regression [MM13, CP15, CD21, WW22], linear programming [CLS21] and more [SWZ19, ${}^{+}$20]. Some of the key matrix sketching techniques we use in this work are subspace embeddings [CW13, NN13, MM13, Coh16, CDDR23], which can be used to construct effective preconditioners for highly over- or under-determined linear systems (e.g., least squares). In particular, our fast least squares solver is an accelerated version of the sketch-to-precondition algorithm by Rokhlin and Tygert [RT08], which, using current subspace embedding techniques, can solve this task in $\tilde{O}(\mathrm{nnz}(\mathbf{A})+d^{\omega})$ time by sketching the tall $n\times d$ matrix $\mathbf{A}$ down to an $\tilde{O}(d)\times d$ size, and then performing a QR/SVD decomposition to produce a preconditioner. When the matrix $\mathbf{A}$ has only $k$ large singular values, then this can be improved to $\tilde{O}(\mathrm{nnz}(\mathbf{A})+d^{2}k^{\omega-2})$, by using randomized low-rank SVD with power iteration [HMT11]. We instead use our stochastic solver to avoid constructing an explicit preconditioner altogether (see Table 1 for comparison). Finally, as mentioned earlier, many randomized low-rank approximation methods mentioned above, such as [${}^{+}$15, CMM17], avoid power iteration and run in linear time. However, these approaches provide a Frobenius norm error guarantee (as opposed to an approximation of all top $k$ singular values), which does not translate into a strong preconditioner for a linear system.
@@ 2 Preliminaries
To construct the preconditioner in our inner iteration for solving the linear system, as well as for obtaining an improved input sparsity time algorithm for least squares, we use the following result which guarantees a $(1+\epsilon)$-subspace embedding with an optimal embedding dimension [CDDR23]. We slightly adapt their result for our setting, so that it holds with a high enough probability to apply a union bound across all iterations of our algorithms.
Lemma 2.7 (Adapted from Theorem 1.4 in [CDDR23]).
@@ 4.3 Approximate Sketch-and-Project
The only part left is to construct a preconditioner $\mathbf{M}\in\mathbb{R}^{\tau\times\tau}$ such that $\tilde{\kappa}=O(1)$, and we construct it by using fast oblivious subspace embedding matrix proposed by [CDDR23]. According to Lemma 2.7, by setting $\phi=O(\tau+\log(1/\delta))$ we have that $\tilde{\kappa}=O(1)$ holds with probability $1-\delta$, since we can set $\epsilon$ to be a small constant. Next we analyze the time complexity: given $\tilde{\mathbf{A}}$, computing $\mathbf{\Theta}\tilde{\mathbf{A}}^{\top}$ takes $O(n\tau\log(\tau/\delta)+\tau^{2}\log^{4}(\tau/\delta))$, computing the SVD of $\mathbf{\Theta}\tilde{\mathbf{A}}^{\top}$ takes $O(\phi\tau^{\omega-1})$, and computing $\mathbf{M}$ takes $O(\tau^{\omega})$. Thus we conclude that the total time complexity is
@@ 6 Replacing DPP with Uniform Sampling
Given a strongly Rayleigh distribution $\mu\in\mathbb{R}^{\tbinom{[m]}{k}}$ and marginal overestimates $q_{i}\geq\Pr\{i\in T\}$ for $i\in[m]$. Let $K:=\sum_{i\in[m]}q_{i}$, there is an algorithm that outputs a sample from a distribution with total variation distance $m^{-O(1)}$ from $\mu$ in time bounded by $O(\log^{3}m)$ calls to $\mathcal{T}_{\mu}(O(K),k)$, where $\mathcal{T}_{\mu}(t,k)$ is the time it takes to produce a sample from $\mu$ conditioned on all elements of the sample being a subset of a fixed set $T$ of size $|T|=t$.

### v2

@@ 1 Introduction
Solving a linear system $\mathbf{A}\mathbf{x}=\mathbf{b}$ is one of the most classical algorithmic tasks, with applications in scientific computing, engineering, data science, optimization, and more. Many classical deterministic algorithms are known for this task, from direct methods such as Gaussian elimination to iterative methods such as Conjugate Gradient. More recently, there has been much success in developing faster algorithms based on randomized techniques such as sampling and sketching for solving linear systems under specific structural conditions, e.g., graph-structured [ST14, KMP12, CKK+18] and other classes of sparse linear systems [XXG12, Gra06], as well as extremely over/under-determined systems [Sar06, Woo14]. However, the benefits of using randomized algorithms for general dense linear systems remain unclear.
@@
The above discussion indicates that solving linear systems with a matrix having only $k=o(n)$ large singular values should be substantially easier than in the worst-case. Understanding this phenomenon is motivated by the real-world prevalence of such linear systems: either due to direct regularization (e.g., $\mathbf{A}=\mathbf{B}+\lambda\mathbf{I}$), as in continuous optimization [BV04], machine learning [ZDW13], and statistics [DW18], or because the input matrix is distorted by isotropic noise (e.g., $\mathbf{A}=\mathbf{B}+\delta\mathbf{N}$, where $\mathbf{N}$ is the noise), which can be caused by measurement error [LW11], data compression [LGW+21], and privacy enforcement [DR14]. Yet, the $\tilde{O}(n^{2}k)$ complexity of CG is highly unsatisfactory, as it only provides an improvement over direct methods for small $k$, and does not lead to a near-linear time algorithm unless $k=\tilde{O}(1)$. This leads to our central question:
@@
Arguably the most popular strategy of accelerating the complexity of linear system solvers such as CG is via preconditioning. For instance, given access to $\mathbf{A}$’s best rank $k$ approximation $\mathbf{A}_{k}$, we can easily construct a preconditioner matrix $\mathbf{M}$ that allows CG to solve our problem in $O(\log(1/\epsilon))$ iterations (since $\mathbf{A}_{k}$ describes the entire ill-conditioned part of $\mathbf{A}$’s spectrum). While computing $\mathbf{A}_{k}$ exactly is as expensive as directly solving the linear system, much recent work has been dedicated to fast randomized algorithms for low-rank approximation via matrix sketching. In particular, we can use power iteration [HMT11] (or Krylov subspace iteration [MM15]) to construct a low-rank approximation $\tilde{\mathbf{A}}_{k}$ with a spectral norm guarantee $\|\tilde{\mathbf{A}}_{k}-\mathbf{A}\|_{2}=O(\sigma_{k+1}(\mathbf{A}))$, sufficient to construct a good preconditioner for our problem. These algorithms require (repeatedly) multiplying matrix $\mathbf{A}$ with an $n\times k$ random matrix, leading to the runtime of $\tilde{O}(n^{2}k^{\omega-2})$. This complexity can be further refined using fast rectangular matrix multiplication [LG12] by replacing the exponent $\omega$ with $\omega(1,1,\log_{n}(k))\leq\omega$. Nevertheless, for large enough $k$, the resulting preconditioning algorithm will not run in near-linear time. [^1] Using current matrix multiplication algorithms, we have $\omega(1,1,\alpha)>2$ for any $\alpha>0.321335$ [WXXZ23]. For simplicity, from now on we will present all time complexities using just the square matrix multiplication exponent $\omega$. Some randomized low-rank approximation algorithms avoid power iteration and run in $\tilde{O}(n^{2}+nk^{\omega-1})$ time [CW13, CEM+15, CMM17], but they only guarantee a bound on the Frobenius norm error $\|\tilde{\mathbf{A}}_{k}-\mathbf{A}\|_{F}$ which is worse by up to a $\sqrt{n}$ factor than the spectral norm bound and thus does not in general lead to a good preconditioner.
@@ 1.1 Main Results
Numerical stability of our methods.
For the sake of simplicity, all our results assume the exact arithmetic (Real-RAM) model. However, we do not foresee any obstacles in establishing the numerical stability of our algorithms in a Word-RAM model with word sizes polylogarithmic in $n$, the condition number of $\mathbf{A}$, and precision $1/\epsilon$. In fact, our convergence analysis of the main algorithmic framework already allows for approximate computations (Section 7), so that an inner solver can be employed to perform them. Thus, the question reduces to the numerical stability of the inner solver. For this, we use preconditioned conjugate gradient (PCG), where the preconditioner is constructed by sketching a tall matrix (i.e., multiplying by a sparse oblivious subspace embedding), and then computing a singular value decomposition of the sketch. While the conjugate gradient method can behave differently in finite precision than in the Real-RAM model, it has been established that it still obtains essentially the same worst-case convergence rate (stated in our Lemma 2.6) in finite precision [Gre89, MMS18]. The preconditioner can be constructed in a stable way using black-box (fast) linear algebra routines [DDH07].
@@ 1.2 Our Techniques
A core inspiration for our algorithms is a broad family of stochastic iterative methods for solving linear systems, which includes randomized Kaczmarz [SV09], coordinate descent [LL10], sketch-and-project [GR15], and many others [LS13]. Unlike CG-type methods which do a pass over the entire matrix $\mathbf{A}$ in each iteration, these approaches randomly sub-sample a small number of rows/columns of $\mathbf{A}$ (or construct a small sketch of $\mathbf{A}$) and use this stochastic information to perform a sublinear iterative update. There has been much success in characterizing the convergence and complexity of stochastic methods in the ultra-small sample regime (i.e., single row/column per iteration [SV09]), but the theoretical analysis of the larger sample regime has proven much more elusive. While there has been some progress in this direction over the last few years, it has come with trade-offs: meaningful characterizations of the convergence behavior have required relying on either constructing sophisticated and expensive sub-sampling distributions (such as determinantal point processes [RK20, MDK20]), or even more expensive dense sketching methods (such as the Gaussian sketch [RN21, DR24]), rendering the entire procedure computationally impractical.
@@ 1.3 Background and Related Work
There has also been much work on accelerating iterative linear system solvers, particularly those based on Krylov subspace methods, such as conjugate gradient / Lanczos. Here, most of the efforts have been concentrated on constructing effective preconditioners, for example via specialized efficient Cholesky factorizations [KLP+16, KS16], or on improving their numerical stability [PV21, Nie22]. However, these approaches have focused on sparse and/or structured systems, e.g., those decomposable into a sum of simple components, particularly Laplacians of undirected graphs [Vai89, ST14, KMP12], and their extensions [CKK+18]. Other examples of structured systems with efficient solvers include Hankel/Toeplitz matrices [KKM79, XXG12] and circulant matrices [Gra06].
@@
Stochastic solvers. The literature on stochastic iterative methods for solving linear systems has its origins in the Kaczmarz method [Kac37], which performs iterative updates while cycling through the rows of the system and projecting the current iterate onto the hyperplane formed by an individual row. A randomized version of this method was introduced and analyzed by Strohmer and Vershynin [SV09], followed up by closely related randomized coordinate descent methods for psd linear systems and least squares by Leventhal and Lewis [LL10]. This triggered much research into stochastic iterative solvers, leading to more general classes of methods that sample blocks of rows instead of a single row [NW13, NT14, MDK20], or use matrix sketching [RN21, DR24] via the Sketch-and-Project framework [GR15]. Despite these efforts, it is still not fully understood when stochastic solvers are faster than deterministic (e.g., Krylov-based) iterative methods. One such result was obtained by Lee and Sidford [LS13], who showed that the convergence rate of an accelerated stochastic coordinate descent solver exhibits better condition number dependence than conjugate gradient for psd linear systems. However, this still leads to poor performance for matrices that have an ill-conditioned top-$k$ part of the spectrum, whereas our algorithms avoid this by using blocks of coordinates (rows or columns) as opposed to one coordinate at a time.
@@
Matrix sketching and sub-sampling. Our techniques are closely related to matrix sketching and sampling techniques commonly used in randomized linear algebra [HMT11, Woo14, DM16, MT20, MDM+23], which is an area that has led to improved randomized approximation algorithms for tasks such as least squares regression [Sar06, RT08, CW13], low-rank approximation [CEM+15, CMM17, LW20], $l_{p}$ regression [MM13, CP15, CD21, WW22], linear programming [CLS21] and more [SWZ19, JKL+20]. Some of the key matrix sketching techniques we use in this work are subspace embeddings [CW13, NN13, MM13, Coh16, CDDR24], which can be used to construct effective preconditioners for highly over- or under-determined linear systems (e.g., least squares). In particular, our fast least squares solver is an accelerated version of the sketch-to-precondition algorithm by Rokhlin and Tygert [RT08], which, using current subspace embedding techniques, can solve this task in $\tilde{O}(\mathrm{nnz}(\mathbf{A})+d^{\omega})$ time by sketching the tall $n\times d$ matrix $\mathbf{A}$ down to an $\tilde{O}(d)\times d$ size, and then performing a QR/SVD decomposition to produce a preconditioner. When the matrix $\mathbf{A}$ has only $k$ large singular values, then this can be improved to $\tilde{O}(\mathrm{nnz}(\mathbf{A})+d^{2}k^{\omega-2})$, by using randomized low-rank SVD with power iteration [HMT11]. We instead use our stochastic solver to avoid constructing an explicit preconditioner altogether (see Table 1 for comparison). Finally, as mentioned earlier, many randomized low-rank approximation methods mentioned above, such as [CEM+15, CMM17], avoid power iteration and run in linear time. However, these approaches provide a Frobenius norm error guarantee (as opposed to an approximation of all top $k$ singular values), which does not translate into a strong preconditioner for a linear system.
@@ 2 Preliminaries
To construct the preconditioner in our inner iteration for solving the linear system, as well as for obtaining an improved input sparsity time algorithm for least squares, we use the following result which guarantees a $(1+\epsilon)$-subspace embedding with an optimal embedding dimension [CDDR24]. We slightly adapt their result for our setting, so that it holds with a high enough probability to apply a union bound across all iterations of our algorithms.
Lemma 2.7 (Adapted from Theorem 1.4 in [CDDR24]).
@@ 4.3 Approximate Sketch-and-Project
The only part left is to construct a preconditioner $\mathbf{M}\in\mathbb{R}^{\tau\times\tau}$ such that $\tilde{\kappa}=O(1)$, and we construct it by using fast oblivious subspace embedding matrix proposed by [CDDR24]. According to Lemma 2.7, by setting $\phi=O(\tau+\log(1/\delta))$ we have that $\tilde{\kappa}=O(1)$ holds with probability $1-\delta$, since we can set $\epsilon$ to be a small constant. Next we analyze the time complexity: given $\tilde{\mathbf{A}}$, computing $\mathbf{\Theta}\tilde{\mathbf{A}}^{\top}$ takes $O(n\tau\log(\tau/\delta)+\tau^{2}\log^{4}(\tau/\delta))$, computing the SVD of $\mathbf{\Theta}\tilde{\mathbf{A}}^{\top}$ takes $O(\phi\tau^{\omega-1})$, and computing $\mathbf{M}$ takes $O(\tau^{\omega})$. Thus we conclude that the total time complexity is
@@ 6 Replacing DPP with Uniform Sampling
Given a strongly Rayleigh distribution $\mu\in\mathbb{R}^{\tbinom{[m]}{k}}$ and marginal overestimates $q_{i}\geq\Pr_{T\sim\mu}\{i\in T\}$ for $i\in[m]$. Let $K:=\sum_{i\in[m]}q_{i}$, there is an algorithm that outputs a sample from a distribution with total variation distance $m^{-O(1)}$ from $\mu$ in time bounded by $O(\log^{3}m)$ calls to $\mathcal{T}_{\mu}(O(K),k)$, where $\mathcal{T}_{\mu}(t,k)$ is the time it takes to produce a sample from $\mu$ conditioned on all elements of the sample being a subset of a fixed set $T$ of size $|T|=t$.